# Secure and Efficient Access Control Framework for Computer-Use Agents via Context Space

### v3

@@ Abstract.
Large language model (LLM)-based computer-use agents represent a convergence of AI and OS capabilities, enabling natural language to control system- and application-level functions. However, due to LLMs’ inherent uncertainty issues, granting agents control over computers poses significant security risks. When agent actions deviate from user intentions, they can cause irreversible consequences. Existing mitigation approaches, such as user confirmation and LLM-based dynamic action validation, still suffer from limitations in usability, security, and performance. To address these challenges, we propose CSAgent, a system-level, static policy-based access control framework for computer-use agents. To bridge the gap between static policy and dynamic context and user intent, CSAgent introduces intent- and context-aware policies, and provides an automated toolchain to assist developers in constructing and refining them. CSAgent enforces these policies through an optimized OS service, ensuring that agent actions can only be executed under specific user intents and contexts. CSAgent supports protecting agents that control computers through diverse interfaces, including API, CLI, and GUI. We implement and evaluate CSAgent, which successfully defends against more than 99.56% of attacks while introducing only 1.99% performance overhead.
@@ 1. Introduction
Recent advances in large language models (LLMs) have paved the way for a new paradigm in human-computer interaction: the computer-use agent (CUA) (sager2025AgentForComputerUse). These agents are capable of autonomously controlling personal computing devices via application programming interfaces (APIs) (xu2024osagent; website:appleaiintent; website:googleaifunc), command-line interfaces (CLIs) (website:anthropicbash; website:langchainbash; website:openaishell), or graphical user interfaces (GUIs) (hong2024cogagent; qin2025uitars; website:openaicua; wen2024autodroid; wang2023enabling; su2025learn; website:anthropiccua), assisting users in daily tasks. Such agents are already being deployed in diverse domains, including PC, smartphones, and in-vehicle systems. By interpreting natural language instructions and orchestrating multi-step workflows across multiple apps, agents significantly enhance user experience and productivity.
@@
However, these agents also expand the attack surface of computer systems and introduce significant security risks, primarily stemming from two factors. First, LLMs are vulnerable to attacks like prompt injection and jailbreaks (greshake2023notwhat; wei2023jailbroken; deng2023masterkeyjailbreak; yu2024dontlistenjailb; shen2024doanythingjailb; liu2024formalizingpromptinj; shi2024optimizationpromptinj), and they have inherent issues like non-determinism and hallucinations, even in well-trained models (huang2025surveyhallucination; sikka2025hallucination). These defects can cause agents to make unpredictable decisions. Second, agents possess permissions comparable to those of users, enabling them to control devices and perform sensitive actions (li2024personal; sager2025AgentForComputerUse; wang2025guiagentsfoundationmodels; zhang2025largelanguagemodelbrainedgui; shi2025trustworthyguiagentssurvey), which violates the principle of least privilege. In combination, these properties can lead agents to execute unintended and unsafe actions, some of which are irreversible, such as deleting data, transferring funds, or unlocking physical devices, which may result in data breaches, financial loss, or even physical harm.
@@
As it is impractical to address these issues through model training alone (wei2023jailbroken; wolf2023fundamental; xu2024hallucination; gong2025safety), commercial agents usually rely on external safeguards to mitigate risks. For example, many agents (website:openaicua; website:anthropiccua; website:vscodeagent) require user confirmations before performing actions. However, this approach burdens the user experience, reduces task efficiency, and thus diminishes overall agent usability. This calls for a more seamless protection mechanism. Inspired by operating systems (OSes), an intuitive approach is to enforce policy-based access control to constrain agent behavior. Drawing from the Contextual Integrity (CI) framework (nissenbaum2009privacy; roesner2012userdrivenac; wijesekera2015androidprem) and Mandatory Access Control (MAC) principles (sandhu1996authentication; wright2002selinux; smalley2013seandroid), we can ensure that security-critical actions are performed only when contextual conditions are satisfied. Recent studies have explored this approach (bagdasarian2024airgapagent; xiang2024guardagent; shi2025progent; tsai2025contextualagent; lee2025verisafe; chen2025shieldagent; yang2025quadsentinelsequentsafetymachinecheckable), and some propose dynamic security policy generation, which enhances security while preserving automation. However, these approaches still suffer from fundamental limitations as they rely on LLM-based dynamic policy generation at agent runtime: unreliable LLMs can produce flawed or incomplete rules, and the inference process incurs significant performance and cost overhead.
@@
First, static policies face difficulties when the security constraints for a function depend on dynamic user intent. For example, file deletion may require path validation for targeted removal but backup verification for cleanup operations. Moreover, users may explicitly specify constraints (e.g., amount limits) in their requests, which also determine how the rules should be defined. Approaches (tsai2025contextualagent; shi2025progent; lee2025verisafe) that generate policies at runtime can easily handle such variations, but they require user instructions, which are not observable at development time. To address this, we propose intent-aware context space, a per-application hierarchical structure that defines and organizes context- and intent-aware policies (§4.2). Context space uses function and intent to index policies, improving the expressiveness to cover a wider range of scenarios. Additionally, we propose an intent prediction method that employs LLMs to predict potential user intents during the policy generation phase.
@@
We implement a CSAgent prototype and conduct extensive evaluation across three benchmarks: AgentBench (liu2023agentbenchevaluatingllmsagents), AgentDojo (agentdojo), and AndroidWorld (rawles2024androidworld), which respectively evaluate CLI-based, API-based, and GUI-based agents. Our experimental results demonstrate that CSAgent successfully blocks more than 99.56% of attacks while introducing only 1.99% average additional latency and 5.42% utility decrease, significantly outperforming existing methods. Additionally, our LLM-based context analyzer identifies 1.93× to 4.12× more GUI elements compared to existing approaches during GUI app analysis, providing a substantially richer semantic knowledge base for automated policy generation.
@@ 2. Background
LLM-based agents are AI systems that use LLMs as their core reasoning engine to perceive environments, make plans, and take actions to accomplish goals (xi2025rise; li2024personal). A key enabler of this paradigm is the function calling capability of LLMs, which allows agents to interface with external systems and perform concrete actions through tools (qin2023toolllm; website:hffunccall; website:oafunccall; du2024anytool). A tool typically refers to a semantic abstraction that encapsulates one or more functions. Computer-use agents (CUAs), as shown in Figure 1, represent a specialized class of LLM agents designed to control computers like humans. Based on how they interface with computers, CUAs can be categorized into three primary approaches: API-based control, CLI-based control, and GUI-based control.
@@
API-based control. One category of tools provides access to the APIs of systems or applications on user devices (website:langchainfs; website:googleaifunc; website:appleaiintent; website:mscopilot). This enables agents to perform high-level actions like document processing, ticket booking, and route planning. API-based control offers significant advantages in reliability and efficiency. Specifically, well-defined APIs provide clear interfaces with predictable behavior, enabling agents to accomplish complex tasks through systematic tool orchestration and deterministic code execution. However, this approach is limited by the predefined tools—agents can only perform actions supported by available APIs. This constraint requires explicit tool development to support new functionalities, limiting the agent’s adaptability to novel user needs and emerging ecosystems.
@@
CLI-based control. Another category of tools leverages command-line interfaces (CLIs) to enable agents to interact with OSes through text commands, providing direct access to system functions and utilities (website:langchainbash; website:anthropicbash; website:openaishell). CLI-based control offers broader functionality compared to API-based approaches, as the extensive ecosystem of CLI tools provides rich functionality without requiring extra API development. However, this flexibility comes with security risks, as command injection vulnerabilities can lead to code execution, privilege escalation, and system compromise (liu2024demystifying). To mitigate these risks, tool providers typically require mechanisms like sandboxing and command filtering (website:anthropicbash; website:openaishell; website:langchainbash).
@@
GUI-based control. A third category of tools enables agents to interact with apps via graphical user interfaces (GUIs) (website:openaicua; website:anthropiccua; wen2024autodroid). GUI is designed only for human users, requiring agents to perceive the screen and manipulate UI elements as a human would. This typically involves capturing screenshots (hong2024cogagent; qin2025uitars) or extracting a structured GUI tree (wen2024autodroid; su2025learn), allowing the agent to understand the current screen, identify actionable elements (such as buttons and input fields), and reason about which actions to take. Finally, the agent performs these actions by calling tools that simulate user interactions, such as clicking or scrolling. GUI-based control offers broad compatibility, allowing agents to interact with any GUI app. This greatly enhances the agent’s adaptability, especially for tasks involving legacy apps that lack API or CLI support. However, this approach still faces several limitations. For example, GUI agents often exhibit low task completion rates due to challenges in GUI comprehension and the complexity of multi-step interactions (rawles2024androidworld; NEURIPS2024osworld). Moreover, they also suffer from poor performance since every action requires LLM inference, resulting in substantial latency and degraded user experience (zhang2025largelanguagemodelbrainedgui; wang2025guiagentsfoundationmodels).
@@ 3.1. Security-usability Trade-offs in CUAs
While CUAs bring convenience to users, they also expand the attack surface of personal devices and introduce security risks. Specifically, due to the inherent uncertainty of LLMs, as well as their susceptibility to attacks such as prompt injection and jailbreaking, agents may perform unexpected actions that deviate from user intentions (li2024personal; sager2025AgentForComputerUse; wang2025guiagentsfoundationmodels), as shown in Figure 1. Many of these actions are irreversible and can cause severe consequences that impact the real world, such as data loss, financial damage, or physical harm.
Although some tools attempt to constrain agents’ capabilities through filtering mechanisms and sandboxing (website:anthropicbash; website:openaishell; website:langchainbash; patil2024goexperspectivesdesignsruntime), these approaches cannot fundamentally solve the problem. Such methods typically either permit or deny specific operations unconditionally, lacking the flexibility to account for the specific context. This creates an inherent dilemma: agents must retain the ability to perform actions to maintain their functionality, yet many of these actions are inherently risky. Prohibiting dangerous operations would limit agent capabilities, while unrestricted access poses unacceptable security risks. This tension between security and functionality is a core challenge in agent design.
A straightforward and widely adopted solution is to request user confirmation before each sensitive action (website:openaicua; website:anthropiccua; website:vscodeagent). However, this approach introduces considerable usability and performance trade-offs. From a usability perspective, frequent confirmation prompts can disrupt the user experience and lead to fatigue, causing users to approve actions without careful consideration (tsai2025contextualagent). From a performance standpoint, the confirmation process introduces non-negligible latency, as each task requires multiple rounds of user interaction and LLM inference cycles. This undermines the primary value proposition of autonomous agents: their ability to complete tasks efficiently without constant human oversight.
@@ 3.2. Rule-based Context Validation for CUAs
Ideally, the execution of a function should be determined based on specific contexts rather than applying unconditional restrictions, aligning with the principles of Contextual Integrity (CI) (nissenbaum2009privacy; roesner2012userdrivenac; wijesekera2015androidprem). Specifically, in our scenario, context refers to the collection of information that characterizes the current execution environment, including user intents, system states, and application-specific parameters. The key insight is that the same action may be safe under certain contextual conditions but dangerous in others. For example, file deletion is safe when the user explicitly specifies the target, but risky without authorization or when targeting critical files. This motivates validation systems that evaluate the appropriateness of agent actions within a given context. Such validation can be performed before (pre-execution) or after (post-execution) each action. While post-execution validation enables more precise anomaly detection by analyzing outcomes, it cannot prevent irreversible harm. Thus, recent studies focus on pre-execution validation. Early work like AirGapAgent (bagdasarian2024airgapagent) applied CI principles, using a separate LLM for data access decisions, but lacked explicit policies, limiting explainability and auditability. GuardAgent (xiang2024guardagent) advanced this by introducing code generation to convert safety guard requests into executable code, offering more structured and deterministic validation than direct LLM decisions.
@@
More recent works have developed policy frameworks that support conditional rules (tsai2025contextualagent; lee2025verisafe; shi2025progent; hu2025agentsentinelendtoendrealtimesecurity). However, these approaches still face the limitation of requiring LLM inference at agent runtime for policy generation, creating the security and efficiency challenges we aim to address. Additionally, policies dynamically generated at different times or devices are likely to vary, making it difficult for policy consistency and preventing the unified improvement of policies through iterative refinement. As shown in Table 1, we analyzed the similarity of policies generated by LLMs multiple times for the same function using apps from AgentDojo (agentdojo). We calculate the policy similarity from both structural and semantic perspectives. Even with a temperature of 0, the overall similarity is only around 0.76, highlighting the uncertainty in policy generation, especially from the semantic perspective.
@@ 3.3. Threat Model and Assumptions
We focus on the risks posed by unreliable LLMs within CUAs. Drawing from the OWASP project (website:owasp), we target the Excessive Agency problem, where agents are granted too much autonomy and perform unintended actions, and Sensitive Information Disclosure, such as inadvertent leakage of private data. These risks can manifest due to various underlying factors, including prompt injection attacks, insecure output handling, and the inherently uncertain nature of current LLMs, such as hallucinations. We assume that the agent framework and user instructions are benign, but the underlying LLM may produce untrustworthy outputs when processing untrusted inputs from external sources like websites or making security-critical decisions. We assume that the software with which the agent interacts (i.e., the OS and legitimate apps) is trustworthy and functions as intended. We trust the agent’s execution framework, tool invocation, and prompt construction operate correctly, and the agent can communicate with our OS service securely. Moreover, context values extracted from these trusted components are also trusted. Our approach does not address attacks targeting LLM training processes or model deployment infrastructure, nor does it cover content safety issues.

### v4

@@ Abstract.
Large language model (LLM)-based computer-use agents represent a convergence of AI and OS capabilities, enabling natural language to control system- and application-level functions. However, due to LLMs’ inherent uncertainty issues, granting agents control over computers poses significant security risks. When agent actions deviate from user intentions, they can cause irreversible consequences. Existing mitigation approaches, such as user confirmation and LLM-based dynamic action validation, still suffer from limitations in usability, security, and performance. To address these challenges, we propose CSAgent, a system-level, static policy-based access control framework for computer-use agents. To bridge the gap between static policy and dynamic context and user intent, CSAgent introduces intent- and context-aware policies, and provides an automated toolchain to assist developers in constructing and refining them. CSAgent enforces these policies through an optimized OS service, ensuring that agent actions can only be executed under specific user intents and contexts. CSAgent supports protecting agents that control computers through diverse interfaces, including API, CLI, and GUI. We implement and evaluate CSAgent, which successfully defends against all attacks in the benchmarks while introducing only 1.99% performance overhead and 5.42% utility decrease.
@@ 1. Introduction
Recent advances in large language models (LLMs) have paved the way for a new paradigm in human-computer interaction: the computer-use agent (CUA) (Sager et al., 2025). These agents are capable of autonomously controlling personal computing devices via application programming interfaces (APIs) (Xu et al., 2024a; Apple, 2025; Google, 2025a), command-line interfaces (CLIs) (Anthropic, 2025a; LangChain, 2025b; OpenAI, 2025c), or graphical user interfaces (GUIs) (Hong et al., 2024; Qin et al., 2025; OpenAI, 2025a; Wen et al., 2024; Wang et al., 2023; Su et al., 2025; Anthropic, 2025b), assisting users in daily tasks. Such agents are already being deployed in diverse domains, including PC, smartphones, and in-vehicle systems. By interpreting natural language instructions and orchestrating multi-step workflows across multiple apps, agents significantly enhance user experience and productivity.
@@
However, these agents significantly expand the attack surface, introducing risks primarily stemming from two factors. First, LLMs are inherently vulnerable to attacks like prompt injection and jailbreaks (Greshake et al., 2023; Wei et al., 2023; Deng et al., 2024; Yu et al., 2024; Shen et al., 2024; Liu et al., 2024c; Shi et al., 2024), and suffer from unpredictability due to hallucinations (Huang et al., 2025; Sikka and Sikka, 2025). Second, agents often possess user-level permissions (Li et al., 2024; Sager et al., 2025; Wang et al., 2025b; Zhang et al., 2025; Shi et al., 2025b), violating the principle of least privilege. Consequently, agents may execute unintended, irreversible actions, such as deleting data, transferring funds, or unlocking devices, resulting in severe financial loss or physical harm.
@@
As it is impractical to address these issues through model training alone (Wei et al., 2023; Wolf et al., 2023; Xu et al., 2024b; Gong et al., 2025), commercial agents usually rely on external safeguards to mitigate risks. For example, many agents (OpenAI, 2025a; Anthropic, 2025b; Code, 2025) require user confirmations before performing actions. However, this approach burdens the user experience, reduces task efficiency, and thus diminishes overall agent usability. This calls for a more seamless protection mechanism. Inspired by operating systems (OSes), an intuitive approach is to enforce policy-based access control to constrain agent behavior. Drawing from the Contextual Integrity (CI) framework (Nissenbaum, 2009; Roesner et al., 2012; Wijesekera et al., 2015) and Mandatory Access Control (MAC) principles (Sandhu and Samarati, 1996; Wright et al., 2002; Smalley and Craig, 2013), we can ensure that security-critical actions are performed only when contextual conditions are satisfied. Recent studies have explored this approach (Bagdasarian et al., 2024; Xiang et al., 2024; Shi et al., 2025a; Tsai and Bagdasarian, 2025; Lee et al., 2025; Chen et al., 2025; Yang et al., 2025), and some propose dynamic security policy generation, which enhances security while preserving automation. However, these approaches still suffer from fundamental limitations as they rely on LLM-based dynamic policy generation at agent runtime: unreliable LLMs can produce flawed or incomplete rules, and the inference process incurs significant performance and cost overhead.
@@
First, static policies face difficulties when the security constraints for a function depend on dynamic user intent. For example, file deletion may require path validation for targeted removal but backup verification for cleanup operations. Moreover, users may explicitly specify constraints (e.g., amount limits) in their requests, which also determine how the rules should be defined. Approaches (Tsai and Bagdasarian, 2025; Shi et al., 2025a; Lee et al., 2025) that generate policies at runtime can easily handle such variations, but they require user instructions, which are not observable at development time. To address this, we propose intent-aware context space, a per-application hierarchical structure that defines and organizes context- and intent-aware policies (§4.2). Context space uses function and intent to index policies, improving the expressiveness to cover a wider range of scenarios. Additionally, we propose an intent prediction method that employs LLMs to predict potential user intents during the policy generation phase.
@@
We implement a CSAgent prototype and conduct extensive evaluation across three benchmarks: AgentBench (Liu et al., 2023), AgentDojo (Debenedetti et al., 2024), and AndroidWorld (Rawles et al., 2024), which respectively evaluate CLI-based, API-based, and GUI-based agents. Our experimental results demonstrate that CSAgent achieves a near-perfect defense rate (blocking 100% of attacks with policy evolution) while introducing only 1.99% average additional latency and 5.42% utility decrease, significantly outperforming existing methods. Additionally, our LLM-based context analyzer identifies 1.93× to 4.12× more GUI elements compared to existing approaches during GUI app analysis, providing a substantially richer semantic knowledge base for automated policy generation.
@@ 2. Background
LLM-based agents are AI systems that use LLMs as their core reasoning engine to perceive environments, make plans, and take actions to accomplish goals (Xi et al., 2025; Li et al., 2024). A key enabler of this paradigm is the function calling capability of LLMs, which allows agents to interface with external systems and perform concrete actions through tools (Qin et al., 2023; Face, 2025; OpenAI, 2025b; Du et al., 2024). A tool typically refers to a semantic abstraction that encapsulates one or more functions. Computer-use agents (CUAs), as shown in Figure 1, represent a specialized class of LLM agents designed to control computers like humans. Based on how they interface with computers, CUAs can be categorized into three primary approaches: API-based control, CLI-based control, and GUI-based control.
@@
API-based control. One category of tools provides access to the APIs of systems or applications on user devices (LangChain, 2025a; Google, 2025a; Apple, 2025; Microsoft, 2025a). This enables agents to perform high-level actions like document processing, ticket booking, and route planning. API-based control offers significant advantages in reliability and efficiency. Specifically, well-defined APIs provide clear interfaces with predictable behavior, enabling agents to accomplish complex tasks through systematic tool orchestration and deterministic code execution. However, this approach is limited by the predefined tools—agents can only perform actions supported by available APIs. This constraint requires explicit tool development to support new functionalities, limiting the agent’s adaptability to novel user needs and emerging ecosystems.
@@
CLI-based control. Another category of tools leverages command-line interfaces (CLIs) to enable agents to interact with OSes through text commands, providing direct access to system functions and utilities (LangChain, 2025b; Anthropic, 2025a; OpenAI, 2025c). CLI-based control offers broader functionality compared to API-based approaches, as the extensive ecosystem of CLI tools provides rich functionality without requiring extra API development. However, this flexibility comes with security risks, as command injection vulnerabilities can lead to code execution, privilege escalation, and system compromise (Liu et al., 2024b). To mitigate these risks, tool providers typically require mechanisms like sandboxing and command filtering (Anthropic, 2025a; OpenAI, 2025c; LangChain, 2025b).
@@
GUI-based control. A third category of tools enables agents to interact with apps via graphical user interfaces (GUIs) (OpenAI, 2025a; Anthropic, 2025b; Wen et al., 2024). GUI is designed only for human users, requiring agents to perceive the screen and manipulate UI elements as a human would. This typically involves capturing screenshots (Hong et al., 2024; Qin et al., 2025) or extracting a structured GUI tree (Wen et al., 2024; Su et al., 2025), allowing the agent to understand the current screen, identify actionable elements (such as buttons and input fields), and reason about which actions to take. Finally, the agent performs these actions by calling tools that simulate user interactions, such as clicking or scrolling. GUI-based control offers broad compatibility, allowing agents to interact with any GUI app. This greatly enhances the agent’s adaptability, especially for tasks involving legacy apps that lack API or CLI support. However, this approach still faces several limitations. For example, GUI agents often exhibit low task completion rates due to challenges in GUI comprehension and the complexity of multi-step interactions (Rawles et al., 2024; Xie et al., 2024). Moreover, they also suffer from poor performance since every action requires LLM inference, resulting in substantial latency and degraded user experience (Zhang et al., 2025; Wang et al., 2025b).
@@ 3.1. Security-usability Trade-offs in CUAs
While CUAs bring convenience to users, they also expand the attack surface of personal devices and introduce security risks. Specifically, due to the inherent uncertainty of LLMs, as well as their susceptibility to attacks such as prompt injection and jailbreaking, agents may perform unexpected actions that deviate from user intentions (Li et al., 2024; Sager et al., 2025; Wang et al., 2025b), as shown in Figure 1. Many of these actions are irreversible and can cause severe consequences that impact the real world, such as data loss, financial damage, or physical harm.
Although some tools attempt to constrain agents’ capabilities through filtering mechanisms and sandboxing (Anthropic, 2025a; OpenAI, 2025c; LangChain, 2025b; Patil et al., 2024b), these approaches cannot fundamentally solve the problem. Such methods typically either permit or deny specific operations unconditionally, lacking the flexibility to account for the specific context. This creates an inherent dilemma: agents must retain the ability to perform actions to maintain their functionality, yet many of these actions are inherently risky. Prohibiting dangerous operations would limit agent capabilities, while unrestricted access poses unacceptable security risks. This tension between security and functionality is a core challenge in agent design.
A straightforward and widely adopted solution is to request user confirmation before each sensitive action (OpenAI, 2025a; Anthropic, 2025b; Code, 2025). However, this approach introduces considerable usability and performance trade-offs. From a usability perspective, frequent confirmation prompts can disrupt the user experience and lead to fatigue, causing users to approve actions without careful consideration (Tsai and Bagdasarian, 2025). From a performance standpoint, the confirmation process introduces non-negligible latency, as each task requires multiple rounds of user interaction and LLM inference cycles. This undermines the primary value proposition of autonomous agents: their ability to complete tasks efficiently without constant human oversight.
@@ 3.2. Rule-based Context Validation for CUAs
Ideally, the execution of a function should be determined based on specific contexts rather than applying unconditional restrictions, aligning with the principles of Contextual Integrity (CI) (Nissenbaum, 2009; Roesner et al., 2012; Wijesekera et al., 2015). Specifically, in our scenario, context refers to the collection of information that characterizes the current execution environment, including user intents, system states, and application-specific parameters. The key insight is that the same action may be safe under certain contextual conditions but dangerous in others. For example, file deletion is safe when the user explicitly specifies the target, but risky without authorization or when targeting critical files. This motivates validation systems that evaluate the appropriateness of agent actions within a given context. Such validation can be performed before (pre-execution) or after (post-execution) each action. While post-execution validation enables more precise anomaly detection by analyzing outcomes, it cannot prevent irreversible harm. Thus, recent studies focus on pre-execution validation. Early work like AirGapAgent (Bagdasarian et al., 2024) applied CI principles, using a separate LLM for data access decisions, but lacked explicit policies, limiting explainability and auditability. GuardAgent (Xiang et al., 2024) advanced this by introducing code generation to convert safety guard requests into executable code, offering more structured and deterministic validation than direct LLM decisions.
@@
More recent works have developed policy frameworks that support conditional rules (Tsai and Bagdasarian, 2025; Lee et al., 2025; Shi et al., 2025a; Hu et al., 2025). However, these approaches still face the limitation of requiring LLM inference at agent runtime for policy generation, creating the security and efficiency challenges we aim to address. Additionally, policies dynamically generated at different times or devices are likely to vary, hindering policy consistency and preventing the unified improvement of policies through iterative refinement. As shown in Table 1, we analyzed the similarity of policies generated by LLMs multiple times for the same function using apps from AgentDojo (Debenedetti et al., 2024). We calculate the policy similarity from both structural and semantic perspectives. Even with a temperature of 0, the overall similarity is only around 0.76, highlighting the uncertainty in policy generation, especially from the semantic perspective.
@@ 3.3. Threat Model and Assumptions
We focus on the risks posed by unreliable LLMs within CUAs. Drawing from the OWASP project (OWASP, 2024), we target the Excessive Agency problem, where agents are granted too much autonomy and perform unintended actions, and Sensitive Information Disclosure, such as inadvertent leakage of private data. These risks can manifest due to various underlying factors, including prompt injection attacks, insecure output handling, and the inherently uncertain nature of current LLMs, such as hallucinations. We assume that the agent framework and user instructions are benign, but the underlying LLM may produce untrustworthy outputs when processing untrusted inputs from external sources like websites or making security-critical decisions. We assume that the software with which the agent interacts (i.e., the OS and legitimate apps) is trustworthy and functions as intended. We trust the agent’s execution framework, tool invocation, and prompt construction operate correctly, and the agent can communicate with our OS service securely. Moreover, context values extracted from these trusted components are also trusted. Our approach does not address attacks targeting LLM training processes or model deployment infrastructure, nor does it cover content safety issues.